PCl5 is 40% dissociated when pressure is 2 atm. It will be 80% dissociated when pressure is approximately.

Grade:11

1 Answers

Trupti baranwal
6 years ago
PCl5(s) = PCl3(l) + Cl2 (g) 0. 0. PxApplying law of mass action ,Kp =Px = 2 ×40/100= 0.8Now , Kp is independent of change in pressure as it only depends only on Temperature changeTherefore, 0.8 = 80/100× P P= 1atm